Sammy's school drama club performed a play. Student tickets cost $2 each, and adult tickets cost $4.50 each. Altogether they sold 80 tickets and made $275. How many adults bought tickets?
Answer by TimothyLamb(4379)

x = students
y = adults
---
x + y = 80
2x + 4.5y = 275
---
put the system of linear equations into standard form
---
x + y = 80
2x + 4.5y = 275
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = students = 34
y = adults = 46
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---